OpenStudy (anonymous):

Solve the system of equations: 2x+5y=16 -5x-2y=2

OpenStudy (paki):

see... Step1: Solve one of the equations for one of the variables. We note that is simplest to solve the second equation for y. 3x+2y=−2x+1=3y Step2: SUBSTITUTE y into first equation. 3x+2(−2x+1)=−2x+1=3y Step3: Solve first equation for x. x=−2x+1=−1y Step4: To find y, substitute −1 for x into second equation. x=y=−1−2⋅(−1)+1 The solution is: x=-1 y=3 You can check the solution using the above calculator.

OpenStudy (anonymous):

2x+5y=16 ------(1 -5x-2y=2-------(2 Multiplying eq (1 by 2 and (2 by 5 and then adding (1 and (2, we find 4x+10y=32 ------(1 -25x-10y=10-------(2 -------------------------- -21 x +0 = 42 x= 42/(-21) x= -2 substituting x=2 in eq (1, we find: 2*(-2)+5y=16 i.e. -4 +5y =16 i.e. 5y = 16 +4 i.e. 5y = 20 i.e. y= 20/5 i.e. y= 4 Hence (-2, 4) are the required solutions of the given system of equations. @lizzieparty

OpenStudy (anonymous):

@paki please chek your solution and answer.